OpenStudy (anonymous):

will fan and medal first to answer. solve the equation: log(x+7)-logx=3

OpenStudy (mathmale):

Please review the basic rules of logs: log a + log b = log ab log a - long b = log (a/b) log a^b = b log a

OpenStudy (mathmale):

Which of those three rules applies to the given question?

OpenStudy (anonymous):

log(x+7)-log(3)=3 log((x+7)/3)=3 {x+7}/3=10^3 x+7=3000 x=2993

OpenStudy (anonymous):

\[\log~m-\log~n=\log \frac{ m }{n }\] \[loga=b,10^{b}=a\]

OpenStudy (anonymous):

@mathmale I have to actually solve for x. That's how to write it as one logarithm

OpenStudy (mathmale):

Yes, of course. Still, one of these three rules of logs applies to the left side of your math problem. Take another look, please.

OpenStudy (anonymous):

Loga-logb=log a/b

OpenStudy (mathmale):

Yes. So, log (x+7) - log (3) = log (what fraction?)

OpenStudy (anonymous):

Why is your second number log(3)?

OpenStudy (anonymous):

it will be log(x+7)-log(x)=log([x+7]/[x]) as loga-logb=log(a//b)

OpenStudy (anonymous):

@rishabhjaiswal225499 As one logarithm yes. I have to solve for x though. every time I try it cancels out

OpenStudy (anonymous):

i have solved it above you may have a look at it x is not getting cancled

OpenStudy (anonymous):

You've put it into one logarithm. I need the value of x. Here are my choices. A. 3.5 B.142.7 C.0.0070 D.0.0707

OpenStudy (mathmale):

Yes. So, log (x+7) - log (3) = log (what fraction? It's important that you answer this. If log a - log b = log a/b, then log (x-7) - log (3) = \[\log \frac{ x+7 }{ 3 }=3.\] Think: How could you solve this for x? What function is the inverse of y=log x? (That's a hint.)

OpenStudy (anonymous):

The second value isn't log(3) its logx

OpenStudy (anonymous):

well than maybe base is e and not 10 if its so than log([x+7]/3)=3 so (x+7)/3=e^3 where e=2.7 so x+7=20.08 x=14.08 i dont know but its crt i am sure

OpenStudy (anonymous):

Here's how I'm working it out and you can tell me what's wrong with it. log(x+7)-logx=3 (x+7)-x=10^3 (x+7)-x=1000 -7 -7 x-x=993 canceled out

OpenStudy (mathmale):

@suckerofmath : Thanks for pointing out my mistake. Can you guide rishabh through the problem solution?

OpenStudy (anonymous):

you know what if loga - logb =x than we cannot write a-b =10^x we have to write log(a/b)=x so a/b=10^x and not a-b=10^x your second step is wrong @suckerofmath

OpenStudy (anonymous):

the 10^3 is how you get rid of the log because you put everything to the power e and it makes the right side of the equation 10^3 @rishabhjaiswal225499

OpenStudy (anonymous):

u can use this \[\log_{a} b=x\] so\[a ^{x}=b\]

OpenStudy (anonymous):

so \[\log_{10} ((x+7)/3)=3 so 10^3=((x+7)/3)\]

OpenStudy (anonymous):

@rishabhjaiswal225499 That still doesn't give the value of x

OpenStudy (mathmale):

\[\log \frac{ x+7 }{ x }=3\]. This is the common log, so the base is 10. Write 10 = 10 and then let log (x+7)/x become the exponent of the first base 10 and 3 the exponent of the second: \[10^{\log \frac{ x+7 }{ 7 }}=10^{3.}\] Can you now solve the problem?

OpenStudy (anonymous):

Now I'm really confused.

OpenStudy (anonymous):